# MedCLIP-SAMv2: Towards Universal Text-Driven Medical Image Segmentation

### v3

@@ 3.2 Zero-shot and Weakly Supervised Medical Image Segmentation
Saliency Map Generation: The embeddings $\mathbf{Z}_{\text{img}}$ and $\mathbf{Z}_{\text{text}}$ are then passed through the Multi-modal Information Bottleneck (M2IB) module, which learns to align the image and text modalities by maximizing the mutual information between them while filtering out irrelevant information between the image embedding and the input image. Specifically, the M2IB module produces a continuous visual saliency map $\mathbf{S}$ for the image representing the importance of each pixel concerning the text prompt. This visual saliency map is produced by optimizing the following objective function:
$\mathbf{S}=MI(\mathbf{Z}_{\text{img}},\mathbf{Z}_{\text{text}};\boldsymbol{\theta})-\beta\times MI(\mathbf{Z}_{\text{img}},\mathbf{I};\boldsymbol{\theta})$ (8)
where $MI$ is the mutual information operation and $\beta$ is a hyperparameter that balances the trade-off between relevance and compression.
K-means Clustering for Initial Segmentation: To generate discrete pixel-wise labels as an initial coarse segmentation, a post-processing step is applied to the saliency map $\mathbf{S}$. This includes applying K-means clustering to divide the pixels in $\mathbf{S}$ into two groups: foreground (region of interests) and background:
$\mathbf{Y}_{\text{kmeans}}=\mathrm{Kmeans}(\mathbf{S};k=2)$ (9)
This separates high-intensity (salient) regions from the background based on pixel intensity in the saliency map to get clusters $C$. Afring, small, disconnected regions arusing $A$ to refine the coarse segmentation Specifically, clusters with an area smaller than $A$ are filtered out as follows:
$\mathbf{Y}_{\text{coarse}}=\{c\in C:\text{Area}(c)>A\}$ (10)
@@
Weakly Supervised Training with Pseudo-labels: To further enhance the segmentation accuracy, the zero-shot segmentation results $\mathbf{Y}_{\text{zero-shot}}$ are then used as pseudo-labels with the input medical images $\mathbf{I}$ to train a segmentation network $\mathbf{M}$ in a weakly supervised manner. Thus, the training data will be $\mathcal{T}$ = $\{(\mathbf{I},\mathbf{Y}_{\text{zero-shot}})\}$. Building on the recent work by Zhao et al. [25], checkpoint ensembling has demonstrated superior effectiveness in uncertainty estimation for medical image segmentation when compared to techniques such as Monte Carlo Dropout and mean-field Bayesian Neural Networks. This finding is particularly relevant in the context of the nnUNet framework [24]. Given a total number of epochs $E$, the training process is divided into $D$ cycles composed of $E_{d}=\frac{E}{D}$ epochs, and during each cycle, we save $G_{d}$ checkpoints of the model. After completing all training cycles, the probabilistic prediction of the final segmentation $\mathbf{Y}_{\text{final}}$ is obtained by averaging the predictions from the $G=D*G_{d}$ total checkpoints saved during the training process providing a Monte-Carlo approximation:
@@ 3.3.1 BiomedCLIP fine-tuning
We employed the public MedPix dataset, which contains various radiological modalities, to fine-tune the BiomedCLIP model [11] with our DHN-NCE loss. The base encoders for images and text were the Vision Transformer (ViT) and PubMedBERT [11], respectively. The MedPix dataset was cleaned by removing special characters, trimming leading and trailing white spaces, and excluding samples with captions shorter than 20 characters. All images were resized to 224 $\times$ 224 pixels and normalized according to the RGB channel means and standard deviations used in the original CLIP model [9]. We performed an 85%-15% split, resulting in 20,292 training images and 3,515 validation images. Fine-tuning was performed with a learning rate of 1E-6, a 50% decay rate, and a batch size of 64.
To test the fine-tuning quality of BiomedCLIP, we assessed the top-1 and top-2 accuracy of matching retrievals for both image-to-text and text-to-image on the ROCO (Radiology Objects in COntext) dataset [41], which contains approximately 7,042 multi-modal medical images covering a wide range of radiological cases. We ran the experiments five times with a batch size of 50, using shuffling to randomize image-text pairs (resulting in 70,420 shuffled examples). In addition, we compared different SOTA loss functions for fine-tuning, including InfoNCE [38], DCL [39] and HN-NCE [42] against our DHN-NCE loss. For a fair comparison, all strategies were trained using the same hyperparameters ($\tau$ = 0.6, learning rate = 1E-6), with the hardness parameters for HN-NCE and DHN-NCE set to $\beta_{1}$ = $\beta_{2}$ = 0.15. We also included baseline results from pre-trained BiomedCLIP [11], PMC-CLIP [43], and CLIP [9].
3.3.2 Image segmentation accuracy
@@
  • Breast Tumor Ultrasound: The Breast Ultrasound Images dataset (BUSI) [44], containing 600 images of benign and malignant tumors for training. Additionally, 50 and 113 images from the UDIAT dataset [45] were used for validation and testing, respectively.
  • Brain Tumor MRI: The Brain Tumor dataset [46], comprising 1,462 T1-weighted MRI scans for training, 1,002 for validation, and 600 for testing.
  • Lung Chest X-ray: The COVID-19 Radiography Database (COVID-QU-Ex) [47, 48] is divided into 16,280 chest X-rays (normal, lung opacity, viral pneumonia, and COVID-19 cases) for training, 1,372 for validation, and 957 for testing.
  • Lung CT: CT scans from [49], consisting of segmentation masks for fibrotic diseased lungs from 107 patients, split into 7,959 slices for training, 3,010 for validation, and 1,800 for testing. The split was done by patient ID to prevent data leakage.
For weakly supervised segmentation, we trained the nnUNet [24] architecture for 600 epochs with 3 cycles for all datasets. The learning rate was initialized to 0.01 and we adopted a cyclical learning rate schedule as described in [25], where the learning rate oscillates between a maximum and minimum value throughout each cycle. This allows the model to escape local optima and explore a wider solution space, leading to more diverse and robust predictions. We saved the last 10 checkpoints in each of the 3 cycles resulting in 30 total model checkpoints. The final segmentation result is averaged from the predictions of these 30 checkpoints and is later thresholded to create a binary mask.
@@
Zero-shot SaLIP 44.3310.12 48.6210.25 47.969.14 50.249.26 63.1411.34 66.4411.58 76.3211.22 78.4611.35 57.9410.48 60.9410.61
SAMAug 56.3910.85 59.2310.92 45.7110.34 48.8111.29 57.1812.12 60.0812.34 44.6110.42 46.4810.57 50.9711.89 53.6512.03
Ours 78.429.32 82.699.50 77.126.82 83.186.70 76.392.14 81.632.15 80.925.51 82.765.44 78.215.70 82.575.55
Weakly Supervised nnUNet 73.7714.48 79.7114.79 77.1612.17 85.2112.60 70.156.40 74.106.59 82.245.12 85.654.70 75.839.54 81.179.67
Ours 78.8712.29 84.5812.19 80.039.91 88.2510.04 80.774.44 84.534.51 88.784.43 91.954.06 82.117.77 87.337.70
One-shot UniverSeg 40.565.14 53.256.22 23.815.45 35.286.49 68.152.21 80.092.16 54.948.21 69.627.59 46.8710.83 59.5611.20
ProtoSAM 48.4410.93 50.2410.84 45.6815.14 51.6915.65 80.751.40 85.111.30 84.509.94 87.629.72 64.8413.01 68.6712.74
Few-shot (K = 4) UniverSeg 47.568.57 54.258.71 53.8210.17 66.409.96 79.252.10 84.801.70 65.6812.02 70.5611.67 61.5811.87 69.0011.31
Self-Prompt-SAM 42.0417.19 44.3017.64 46.4315.25 50.2915.83 67.972.89 71.632.83 81.503.84 83.403.77 59.4913.10 62.4113.06
Few-shot (K = 16) UniverSeg 66.368.57 72.228.30 62.827.97 72.767.94 83.441.54 87.731.24 86.492.49 89.961.94 74.787.39 80.676.25
Self-Prompt-SAM 62.3616.38 66.0116.92 52.5515.29 57.0715.93 82.492.50 86.492.45 83.663.90 85.493.84 70.2711.84 73.7711.69
Fully Supervised nnUNet 82.4710.49 88.3210.77 87.746.28 95.106.28 98.720.65 99.510.41 97.102.74 99.182.13 84.635.04 90.424.80
nnUNet Ensemble 84.7210.97 90.8511.26 88.825.93 95.845.54 99.142.50 99.821.93 98.124.09 99.654.03 85.435.87 91.745.71
@@ 3.3.3 Experimental setup and metrics
We performed a comprehensive comparison of segmentation quality using the initial labels derived from post-processed M2IB results, zero-shot pseudo-masks, and weakly supervised outputs on the specified testing datasets. Our zero-shot method was benchmarked against SOTA zero-shot (SaLIP [50] and SAMAug [51]) and few-shot (UniverSeg [52], ProtoSAM [53], and Self-Prompt-SAM [54]) segmentation methods. Additionally, we compare our weakly supervised method with nnUNet [24] trained on pseudo-labels without checkpoint ensembling. As part of the ablation studies for zero-shot segmentation, we examined: 1) the impact of fine-tuning BiomedCLIP and the choice of explainable AI (XAI) technique for saliency map generation, 2) the influence of different text prompts on overall segmentation performance, 3) the contribution of each model component to the final performance, and 4) the selection of SAM pre-trained models with various visual prompting strategies. These ablation studies were conducted on the test sets of all four datasets mentioned. In all experiments, Dice-Sørensen Coefficient (DSC) and Normalized Surface Distance (NSD) were used as evaluation metrics. Paired-sample t-tests were also conducted to validate the observed trends, with a p-value of less than 0.05 indicating statistical significance.
@@ 4.1 Comparison with SOTA Methods
Table 1 shows a comparison of the proposed methods with different SOTA techniques. In general, our method demonstrated superior performance in both zero-shot and weakly supervised settings. In the weakly supervised scenario, our method achieved an average DSC of 82.11 and NSD of 87.33, significantly outperforming all zero-shot and few-shot SOTA methods across various imaging modalities/tasks ($p$ $<$ 0.05), except Lung X-ray. However, the fully supervised methods still offer higher accuracy than those using limited resources. Notably, our zero-shot approach achieved an average DSC of 78.21 and NSD of 82.57, even surpassing weakly supervised nnUNet trained on pseudo-labels without checkpoint ensembling on average.
4.2 Cross-modal retrieval accuracy
The accuracy of cross-modal retrieval (text-to-image and image-to-text) for the ROCO dataset [41] is shown in Table 2 across different losses for fine-tuning BiomedCLIP, with three pre-trained CLIP models as baselines. It can be seen that domain-specific pre-trained models performed better than CLIP, with the larger-scale pretraining offering better retrieval accuracy while the pre-trained BiomedCLIP demonstrating the highest retrieval accuracy among all pre-trained models. Fine-tuning BiomedCLIP further enhanced its performance. Specifically, BiomedCLIP fine-tuned with DHN-NCE reached 84.70% top-1 and 94.73% top-2 in image-to-text retrieval, and 85.99% top-1 and 95.17% top-2 in text-to-image retrieval, significantly outperforming other loss functions and the baseline models ($p<$ 0.01). Additionally, the benefit of fine-tuning BiomedCLIP with our DHN-NCE loss is further validated with improved segmentation quality across different tasks and image modalities in Table 4 and Table 5.
Model Version $image\rightarrow text$ (%) $text\rightarrow image$ (%)
Top-1 Top-2 Top-1 Top-2
CLIP [9] Pre-trained 26.680.30 41.800.19 26.170.20 41.130.20
PMC-CLIP [43] Pre-trained 75.470.37 87.460.11 76.780.11 88.350.19
BiomedCLIP [11] Pre-trained 81.830.20 92.790.13 81.360.48 92.270.14
InfoNCE [38] 84.210.35 94.470.19 85.730.19 94.990.16
DCL [39] 84.440.37 94.680.19 85.890.16 95.090.19
HN-NCE [42] 84.330.35 94.600.19 85.800.17 95.100.19
DHN-NCE 84.700.33 94.730.16 85.990.19 95.170.19

### v4

@@ 3.2 Zero-shot Medical Image Segmentation
LLM Prompt Generation: Since BiomedCLIP utilizes text captions from PubMed to pre-train its text encoder (i.e., PubMedBERT), we utilize LLMs like GPT-4 (Achiam et al., 2023) to generate sophisticated text prompts that can guide the model to localize certain salient regions. Specifically, we generate descriptive textual prompts that can guide the model to pay attention to salient features in the medical scan as follows:
Give a sentence caption that describes unique visual features of [TARGET] in [MODALITY]
We can engineer different prompt configurations ranging from generic to class-specific context captions and we study the effect of these different styles in Section 4.3.1.
Saliency Map Generation: The embeddings $\mathbf{Z}_{\text{img}}$ and $\mathbf{Z}_{\text{text}}$ are then passed through the Multi-modal Information Bottleneck (M2IB) module (Wang et al., 2024), which learns to align the image and text modalities by maximizing the mutual information between the input image $\mathbf{I}$ and a good representative text prompt $\mathbf{T}$ while filtering out irrelevant information between the image embedding and the input image. By doing so, the process bridges the semantic gap between encoded visual and textual features ensuring that embeddings emphasize the content that is jointly relevant across image and text. Specifically, the M2IB module introduces a stochastic information bottleneck $\boldsymbol{\lambda}_{S}$in $\mathbb{R}^{H\times W}$ such that $0\leq\boldsymbol{\lambda}_{S}\leq 1$ where $H$,$W$ are the respective height and weight of the input image $\mathbf{I}$. This produces a continuous visual saliency map for the image representing the importance of each pixel concerning the text prompt. This visual saliency map is produced by optimizing the following objective function:
$\boldsymbol{\lambda_{S}}=MI(\mathbf{Z}_{\text{img}},\mathbf{Z}_{\text{text}};\boldsymbol{\theta})-\gamma\times MI(\mathbf{Z}_{\text{img}},\mathbf{I};\boldsymbol{\theta})$ (16)
where $MI$ is the mutual information operation and $\gamma$ is a hyperparameter that balances the trade-off between relevance and compression.
Post-processing for Initial Segmentation: To obtain a discrete pixel-wise segmentation, we apply Otsu’s thresholding (Otsu, 1979) to the saliency map $\boldsymbol{\lambda_{S}}$, automatically determining an optimal threshold $\eta^{*}$ that separates foreground (regions of interest) from background by minimizing intra-class variance. The binarized segmentation is then given by:
$\mathbf{Y}_{\text{otsu}}=\begin{cases}1,&\boldsymbol{\lambda}_{S}(x,y)\geq\eta^{*}\\ 0,&\boldsymbol{\lambda}_{S}(x,y)<\eta^{*}\end{cases}$ (17)
After thresholding, small, disconnected contours may still exist. To refine the segmentation and ensure robust results, we perform Connected Component Analysis on the identified contours $C$. For each connected component $c\in C$, we compute a confidence score based on the saliency map $\boldsymbol{\lambda_{S}}$. The confidence of a connected component is derived as follows:
$\text{Confidence}(c)=\frac{\sum\limits_{i\in c}p_{i}\cdot\hat{y}_{i}}{\sum\limits_{i\in c}\hat{y}_{i}},$ (18)
where $p_{i}$ is the probability that pixel $i$ belongs to the foreground class, and $\hat{y}_{i}$ is the predicted binary label for pixel $i$. Using this confidence score, we identify the most confident connected components to form the final coarse segmentation:
$\mathbf{Y}_{\text{coarse}}=\{c\in C:\text{Confidence}(c)>\eta_{c}\},$ (19)
where $\eta_{c}$ is a confidence threshold. This process refines the initial segmentation by removing unreliable regions and retaining only high-confidence contours.
@@
3.3 Uncertainty-Aware Weakly Supervised Segmentation
To further enhance the segmentation accuracy, the zero-shot segmentation results $\mathbf{Y}_{\text{zero-shot}}$ are then used as pseudo-labels with the input medical images $\mathbf{I}$ to train a segmentation network $\mathbf{M}$ in a weakly supervised manner. Thus, the training data will be $\mathcal{T}$ = $\{(\mathbf{I},\mathbf{Y}_{\text{zero-shot}})\}$. Building on the recent work by Zhao et al. (Zhao et al., 2022), checkpoint ensembling has demonstrated superior effectiveness in uncertainty estimation for medical image segmentation when compared to techniques such as Monte Carlo Dropout and mean-field Bayesian Neural Networks. This finding is particularly relevant in the context of the nnUNet framework (Isensee et al., 2021). Given a total number of epochs $E$, the training process is divided into $D$ cycles composed of $E_{d}=\frac{E}{D}$ epochs, and during each cycle, we save $G_{d}$ checkpoints of the model. Importantly, this checkpoint strategy adds no delays to the training process, as it involves saving alternate checkpoints of the same model rather than training separate models. After completing all training cycles, the probabilistic prediction of the final segmentation $\mathbf{Y}_{\text{final}}$ is obtained by averaging the predictions from the $G=D*G_{d}$ total checkpoints saved during the training process providing a Monte-Carlo-like approximation:
@@ 3.4.1 BiomedCLIP fine-tuning
We employed the public MedPix (Siragusa et al., 2024) dataset, which contains various radiological modalities, to fine-tune the BiomedCLIP model (Zhang et al., 2023) with our DHN-NCE loss. The base encoders for images and text were the Vision Transformer (ViT) and PubMedBERT (Zhang et al., 2023), respectively. The MedPix dataset was cleaned by removing special characters, trimming leading and trailing white spaces, and excluding samples with captions shorter than 20 characters. All images were resized to 224 $\times$ 224 pixels and normalized according to the RGB channel means and standard deviations used in the original CLIP model (Radford et al., 2021). We performed an 85%-15% split, resulting in 20,292 training images and 3,515 validation images. Fine-tuning was performed with a learning rate of 1E-6, a 50% decay rate, and a batch size of 64.
To validate the fine-tuning quality of BiomedCLIP, we assessed the top-1 and top-2 accuracy of matching retrievals for both image-to-text and text-to-image on the ROCO (Radiology Objects in COntext) dataset (Pelka et al., 2018), which contains approximately 7,042 multi-modal medical images covering a wide range of radiological cases. We ran the experiments five times with a batch size of 50, using shuffling to randomize image-text pairs (resulting in 70,420 shuffled examples). In addition, we compared different SOTA loss functions for fine-tuning, including InfoNCE (Oord et al., 2018), DCL (Yeh et al., 2022) and HN-NCE (Radenovic et al., 2023) against our DHN-NCE loss. For a fair comparison, all strategies were trained using the same hyperparameters ($\tau$ = 0.6, learning rate = 1E-6), with the hardness parameters for HN-NCE and DHN-NCE set to $\beta_{1}$ = $\beta_{2}$ = 0.15. As a reference, we also included baseline results from pre-trained BiomedCLIP (Zhang et al., 2023), PMC-CLIP (Lin et al., 2023a), and CLIP (Radford et al., 2021).
3.4.2 Datasets
@@
  • 1.
    Breast Tumor Ultrasound: The Breast Ultrasound Images dataset (BUSI) (Al-Dhabyani et al., 2020), containing 600 images of benign and malignant tumors for training. Additionally, 50 and 113 images from the UDIAT dataset (Byra et al., 2020) were used for validation and testing, respectively.
  • 2.
    Brain Tumor MRI: The Brain Tumor dataset (Cheng, 2017), comprising 1,462 T1-weighted MRI scans for training, 1,002 for validation, and 600 for testing.
  • 3.
    Lung Chest X-ray: The COVID-19 Radiography Database (COVID-QU-Ex) (Chowdhury et al., 2020; Rahman et al., 2021) is divided into 16,280 chest X-rays (normal, lung opacity, viral pneumonia, and COVID-19 cases) for training, 1,372 for validation, and 957 for testing.
  • 4.
    Lung CT: CT scans from (Konya, 2020), consisting of segmentation masks for fibrotic diseased lungs from 107 patients, split into 7,959 slices for training, 3,010 for validation, and 1,800 for testing. The split was done by patient ID to prevent data leakage.
@@
Zero-shot SaLIP 44.3310.12 48.6210.25 47.969.14 50.249.26 63.1411.34 66.4411.58 76.3211.22 78.4611.35 57.9410.49 60.9410.65
SAMAug 56.3910.85 59.2310.92 45.7110.34 48.8111.29 57.1812.12 60.0812.34 44.6110.42 46.4810.57 50.9710.96 53.6511.30
MedCLIP-SAM 67.828.26 69.129.12 66.725.27 68.016.16 64.499.09 65.8910.44 59.149.52 60.479.98 64.548.20 66.109.08
Ours 77.769.52 81.119.89 76.527.06 82.237.13 75.793.44 80.883.52 80.385.81 82.035.94 77.616.82 81.567.00
Weakly Supervised nnUNet 73.7714.48 79.7114.79 77.1612.17 85.2112.60 70.156.40 74.106.59 82.245.12 85.654.70 75.8310.31 81.1710.52
MedCLIP-SAM 58.625.66 60.945.87 58.808.63 61.778.64 86.078.61 88.658.09 80.128.38 83.738.29 70.907.92 73.777.80
Ours 78.8712.29 84.5812.19 80.039.91 88.2510.04 80.774.44 84.534.51 88.784.43 91.954.06 82.118.49 87.338.46
One-shot UniverSeg 40.565.14 53.256.22 23.815.45 35.286.49 68.152.21 80.092.16 54.948.21 69.627.59 46.875.67 59.565.98
ProtoSAM 48.4410.93 50.2410.84 45.6815.14 51.6915.65 80.751.40 85.111.30 84.509.94 87.629.72 64.8410.60 68.6710.71
Few-shot (K = 4) UniverSeg 47.568.57 54.258.71 53.8210.17 66.409.96 79.252.10 84.801.70 65.6812.02 70.5611.67 61.589.02 69.008.86
Self-Prompt-SAM 42.0417.19 44.3017.64 46.4315.25 50.2915.83 67.972.89 71.632.83 81.503.84 83.403.77 59.4911.74 62.4112.08
Few-shot (K = 16) UniverSeg 66.368.57 72.228.30 62.827.97 72.767.94 83.441.54 87.731.24 86.492.49 89.961.94 74.786.03 80.675.86
Self-Prompt-SAM 62.3616.38 66.0116.92 52.5515.29 57.0715.93 82.492.50 86.492.45 83.663.90 85.493.84 70.2711.44 73.7711.84
Fully Supervised nnUNet 82.4710.49 88.3210.77 87.746.28 95.106.28 98.720.65 99.510.41 97.102.74 99.182.13 84.636.27 90.426.33
nnUNet Ensemble 84.7210.97 90.8511.26 88.825.93 95.845.54 99.142.50 99.821.93 98.124.09 99.654.03 85.436.68 91.746.66
@@
Model Version $image\rightarrow text$ (%) $text\rightarrow image$ (%)
Top-1 Top-2 Top-1 Top-2
CLIP (Radford et al., 2021) Pre-trained 26.680.30 41.800.19 26.170.20 41.130.20
PMC-CLIP (Lin et al., 2023a) Pre-trained 75.470.37 87.460.11 76.780.11 88.350.19
BiomedCLIP (Zhang et al., 2023) Pre-trained 81.830.20 92.790.13 81.360.48 92.270.14
InfoNCE (Oord et al., 2018) 84.210.35 94.470.19 85.730.19 94.990.16
DCL (Yeh et al., 2022) 84.440.37 94.680.19 85.890.16 95.090.19
HN-NCE (Radenovic et al., 2023) 84.330.35 94.600.19 85.800.17 95.100.19
DHN-NCE (ours) 84.700.33 94.730.16 85.990.19 95.170.19
Prompt Breast Ultrasound Brain MRI Lung X-ray Lung CT
P0 63.7915.12 67.8915.08 70.987.61 76.427.63 75.793.44 80.883.52 69.895.14 71.834.98
P1 67.6614.35 71.5614.78 37.1910.98 39.7711.63 69.724.65 73.524.83 - -
P2 69.0412.45 73.3312.97 71.187.16 77.197.14 63.914.73 67.635.13 80.385.81 82.035.94
P3 77.769.52 81.119.89 76.527.06 82.237.13 63.924.88 67.734.96 - -
P4 67.6516.54 71.0216.89 69.238.41 74.328.59 68.954.91 72.314.95 75.844.88 77.564.97
P5 65.1817.51 68.7517.93 69.817.86 75.017.97 68.444.63 72.094.81 - -
Table 3: Effect of different text prompt templates on the segmentation performance (%, meanstd)
@@ 3.4.3 Experimental setup and metrics
We performed a comprehensive comparison of segmentation quality using the initial labels derived from post-processed M2IB results, zero-shot pseudo-masks, and weakly supervised outputs on the specified testing datasets. Our zero-shot method was benchmarked against SOTA zero-shot segmentation methods, such as SaLIP (Aleem et al., 2024) and SAMAug (Dai et al., 2024) and few-shot approaches, such as UniverSeg (Butoi et al., 2023), ProtoSAM (Ayzenberg et al., 2024), and Self-Prompt-SAM (Wu et al., 2023b). Additionally, we compare our weakly supervised method with nnUNet (Isensee et al., 2021) trained on pseudo-labels without checkpoint ensembling. For weakly supervised segmentation, we trained the nnUNet (Isensee et al., 2021) architecture for 600 epochs with 3 cycles for all datasets. The learning rate was initialized to 0.01 and we adopted a cyclical learning rate schedule as described in (Zhao et al., 2022), where the learning rate oscillates between a maximum and minimum value throughout each cycle. This allows the model to escape local optima and explore a wider solution space, leading to more diverse and robust predictions. We saved the last 10 checkpoints in each of the 3 cycles resulting in 30 total model checkpoints. The final segmentation result is averaged from the predictions of these 30 checkpoints and is later thresholded to create a binary mask.
As part of the ablation studies for zero-shot segmentation, we examined: 1) the impact of fine-tuning BiomedCLIP and the choice of explainable AI (XAI) technique for saliency map generation, 2) the influence of different text prompts on overall segmentation performance, 3) the contribution of each model component to the final performance, and 4) the selection of SAM pre-trained models with various visual prompting strategies. These ablation studies were conducted on the test sets of all four datasets mentioned.
In all experiments, Dice-Sørensen Coefficient (DSC) and Normalized Surface Distance (NSD) were used as evaluation metrics. Paired-sample t-tests were also conducted to validate the observed trends, with a p-value of less than 0.05 indicating statistical significance.
@@ 4.1 Comparison with SOTA Methods
Table 1 shows a comparison of the proposed MedSAM-CLIPv2 with different SOTA techniques. Compared to the original MedCLIP-SAM, our approach significantly improved the average DSC from 64.54% to 77.61% and NSD from 66.10% to 81.56% in the zero-shot setting. Similarly, in the weakly supervised scenario, the average DSC increased from 70.90% to 82.11% and NSD from 73.77% to 87.33%, even surpassing weakly supervised nnUNet trained on pseudo-labels without checkpoint ensembling on average. Overall, our method significantly outperformed all zero-shot and few-shot SOTA methods across various imaging modalities/tasks ($p$ $<$ 0.05), except for Lung X-ray. However, the fully supervised methods still offer higher accuracy than those using limited resources.
4.2 Effectiveness of DHN-NCE
The accuracy of cross-modal retrieval (text-to-image and image-to-text) for the ROCO dataset (Pelka et al., 2018) is shown in Table 2 across different losses for fine-tuning BiomedCLIP, with three pre-trained CLIP models as baselines. It can be seen that domain-specific pre-trained models performed better than CLIP, with the larger-scale pretraining offering better retrieval accuracy while the pre-trained BiomedCLIP demonstrating the highest retrieval accuracy among all pre-trained models. Fine-tuning BiomedCLIP further enhanced its performance. Specifically, BiomedCLIP fine-tuned with DHN-NCE reached 84.70% top-1 and 94.73% top-2 in image-to-text retrieval, and 85.99% top-1 and 95.17% top-2 in text-to-image retrieval, significantly outperforming other loss functions and the baseline models ($p<$ 0.01). Additionally, the benefit of fine-tuning BiomedCLIP with our DHN-NCE loss is further validated with improved segmentation quality across different tasks and image modalities in Table 4 and Table 5.